Cash 68,000 22,000 Accounts Receivable 88,000 76,000 Inventory 167,000 189,000 Land 80,000 100,000 Equipment 260,000 200,000...

Cash

68,000

22,000

Accounts Receivable

88,000

76,000

Inventory

167,000

189,000

Land

80,000

100,000

Equipment

260,000

200,000

Accumulated Depreciation

-66,000

-32,000

Total

597,000

555,000

Liabilities and Stockholders' Equity

Accounts Payable (for purchases of inventory)

39,000

43,000

Bonds Payable

150,000

200,000

Common Stock

216,000

174,000

Retained Earnings

192,000

138,000

Total

597,000

555,000

Additional Information

1.  Net income for 2022 was $93,000

2.  Depreciation expense was $34,000

3.  Cash dividends of $39,000 were declared and paid.

4.  Bonds payable amounting to $50,000 were redeemed for cash $50,000

5. Common stock was issued for $42,000 cash

6.  No equipment was sold during 2022.

7.  Land was sold for its book value.

B.  If you prepared Cash Flow from Operations (CFO) using the direct method, you would need to indicate cash inflows and outflows on a line-by-line basis.  

Provide the following amounts:

a.  Cash inflows from customers

b.  Cash payments for inventory
Answer #1

Under Direct method actual cash inflow and outflow from Customers are Consider instead of Adjusting income for cash flow .

Cash Flow from Operating Activity

Cah Inflow From Customer
Total Sales (Note 3) 149000
Less Credit Sales (Note 3) (12000)
Cash flow from Customer 137000
Cash Payment for Inventory ( Note 4) (4000)
Net Cash flow from Operating Activities 133000

Please Check note 3 for Credit Sales

Under Operating Activites , Cash floows relating to day to day business activities , such as sales , purchase , Expenses , which are part of income statement are comes under Operating Activities

1) Though Depreciation is part of Operating activites but it is just amortisation of Fixed Assets expenes , and it does not cause any cash outflow , where on purchase of Fixed assts it is comes under Investing Activities as Investment on Fixed Assets are Part of Investment Activites .

2) Dividend are for Sharehokder , Funds are raise from issue of Share , which are part of Financing activity , hence payment of Dividend are part of Financing Activity

3) Net Income = Total Sales - Cost of Goods Sold - Depreciation

Total Sales to Customer = Net Income + Depreciation + Cost of Goods Sold (Reduction in Inventory )

=93000+34000+22000

=149000

Total Sales = Cash sales + Credit Sales

Cash Sales = Total Sales - Credit Sales

Cash Sales = 149000-12000( Increse In Account Receivables )

=137000

4) CashPayment to Inventory = Decrease in Account Payable from Last yeat to Current

=43000-39000

=4000